# Tail Index Estimation for Discrete Heavy-Tailed Distributions with Application to Statistical Inference for Regular Markov Chains

### v2

@@ 1 Introduction
where $L:\mathbb{R}_{+}\to\mathbb{R}$ is a slowly varying function, i.e. such that $L(\lambda z)/L(z)\rightarrow+1$ as $z\rightarrow+\infty$ for any $\lambda>0$, see [7]. Such discrete power law probability distributions, also referred to as generalized Zipf’s laws sometimes, are often used to model the distribution of discrete data exhibiting a specific rank-frequency relationship, namely when the logarithm of the frequency and that of the rank order are nearly proportional. Such a phenomenon has been empirically observed in many ranking systems: in quantitative linguistics (i.e. when analysing word frequency law in natural language, see e.g. [28]) in the first place, as well as in a very wide variety of situations, too numerous to be exhaustively listed here. One may refer to [35], [25] or [37] among many others. In this paper, we first consider the issue of estimating the parameter $\beta$ involved in (1) (supposedly unknown, like the function $L$) in the classic (asymptotic) i.i.d. statistical setting, i.e. based on an increasing number $n\geq 1$ of independent copies $S_{1},\;\ldots,\;S_{n}$ of the generic r.v. $S$. Statistical inference for discrete heavy-tailed distributions has not received much attention in the literature and still poses methodological problems that this article seeks to resolve. Most of the very few dedicated methods documented either deal with very specific cases as in e.g. [16], [27] and [11] or else consist in empirically applying techniques originally designed for continuous heavy-tailed distributions to discrete data after first adding independent uniform noise [36].
@@
The vast majority of the regular variation index estimators proposed in the literature, Hill or Pickand estimators in particular [20, 33], are based on order statistics, which causes obvious difficulties in the discrete case because of the possible occurrence of many ties. Indeed, in that case, many spacings are equal to 00 which causes these estimators to behave erroneously. This is stressed [27] (see their figures 1 and 2). More importantly, [27] constructs an example of a family of discrete value heavy-tailed distributions that do not satisfy any second-order condition so that the asymptotic normality of the Hill estimator cannot be proved directly. In contrast, the estimator under study here is based on the analysis of the probability of exponentially separated tail events. It simply rests on the fact that, as can be immediately deduced from (1), we have $\ln(p_{k})-\ln(p_{k+1})=\beta+\ln(L(e^{k})/L(e^{k+1}))$, where $\ln(x)$ denotes the natural logarithm of any real number $x>0$ and $p_{l}=\mathbb{P}(S>e^{l})$ for all $l\in\mathbb{N}$, and that $L(e^{k+1})/L(e^{k})$ is expected to be very close to $1$ for $k\in\mathbb{N}$ chosen sufficiently large. A natural (plug-in) inference technique can be then devised by replacing the tail probabilities $p_{l}$ with their empirical versions $\widehat{p}^{(n)}_{l}=(1/n)\sum_{i=1}^{n}\mathbb{I}\{S_{i}>e^{l}\}$ for $l\in\mathbb{N}$, where $\mathbb{I}\{\mathcal{A}\}$ means the indicator function of any event $\mathcal{A}$. This yields the estimator
@@
Although estimation of the parameter $\beta>0$ in (1) in the discrete i.i.d. setting is an important issue in itself, the present paper also finds its motivation in the problem of recovering statistically the regularity index $\beta\in(0,1)$ of a regenerative regular $\beta$-null-recurrent Markov chain $X=(X_{n})_{n\in\mathbb{N}}$, based on the observation of a finite sample path $X_{1},\;\ldots,\;X_{n}$ with $n\geq 1$. As explained in [8] (see also [9]), for regular Markov chains, the regularity index $\beta\in(0,1]$ controls the (sublinear) rate at which the number of visits to any given Harris set increases with observation time $n$, no matter the initial distribution. Recall that this kind of Markov chain is strongly non-stationary and are not mixing in any sense: random walks and their variations, for instance Bessel random walks (see [2], are typical examples. Such processes appear also naturally in many applications: for population dynamics or stochastic growth models in demography, see for instance [1] and the references therein; for branching processes, Galton-Watson processes with immigration (or population dependent) also exhibit a fat tail behaviour of the time returns of the chain, see [32, 38, 22].
@@
In the regenerative case (i.e. when the chain $X$ possesses an accessible atom, a Harris set on which the transition probability is constant), the distribution of the regenerative time, the return time to the atom, is a discrete generalized Pareto (1) and the parameter $\beta$ is its tail index. Due to the non-standard behaviour of traditional estimators in this context, statistical inference for null-recurrent Markov chains is very poorly documented in the literature [24, 23, 26, 18] and, to the best of our knowledge, estimation of the key quantity $\beta$ has not received much attention. It is also the goal of this article to extend the use of the estimator (2) to the case where the $S_{i}$’s are the successive durations between the consecutive regeneration times up to time $n$. The main difficulty naturally arises from the fact that the number $1+N_{n}\geq 0$ of regeneration times (and thus the number of durations) is now random, and the variables $S_{1},\ldots,S_{N_{n}}$ are not independent anymore when $N_{n}\geq 1$ (in particular, their sum is less than $n$ by construction). We show that the estimator remains strongly consistent in this scenario. For illustration purposes, numerical experiments have been carried out, providing empirical evidence of the relevance of the estimation method promoted. Extension to the general case of (pseudo-regenerative) null-recurrent chains is also discussed, the difficulties inherent in applying the methodology originally proposed in [6] in the positive recurrent case to mimic regenerative Nummelin extensions [31] being explained at length.
@@
The paper is organized as follows. A thorough analysis of the behaviour of the estimator (2) in the i.i.d. case, illustrated by numerical experiments, is first carried out in section 2. The asymptotic results thus established are next extended in section 3 to the regenerative regular Markovian setup, when the estimator is computed based on a single finite-length trajectory of the atomic chain. Experimental results are also displayed and the main barrier to the extension of the methodology promoted to general (i.e. pseudo-regenerative) regular null-recurrent chains is also discussed therein. Technical proofs are deferred to the Appendix section.
@@ 2.1 Main Results - Confidence Bounds and Limit Theorems
Suppose that, as $n\rightarrow+\infty$, we have $k_{n}\to+\infty$ so that $(\ln n)\exp(k_{n}\beta)/n=o(L(\exp(k_{n}))$. Then, we have:
@@
In particular, as stated below, strong consistency is guaranteed when $k_{n}$ is of logarithmic order. Notice that this is in an improvement with respect to Hill’s estimator, as in that case, one must choose $k_{n}=o(\ln\ln n)$ in order to guarantee strong consistency, see [13].
@@
Of course, the condition (5) on $k_{n}$ can be hardly checked in practice. This is a classic issue in tail estimation and in the statistical analysis of extreme values more generally. The choice of the hyperparameter $k$ somehow rules the (asymptotic) bias-variance trade-off: the estimator (2) is expected to be of large variance when $k$ is large and to have a large bias if $k$ is too small. As depicted in Fig. 1(a), to choose $k$, one may use the same approach as that originally proposed for the Hill estimator (see e.g. [34]), which consists in plotting the values of (2) for a range of values of $k$ and choosing $k$ in a region where a certain degree of stability is exhibited.
@@
Figure 1: (a) Behaviour of $\widehat{\beta}_{n}(k)$ for different values of $k$, to estimate the parameter $\beta=0.8$ based on a dataset of $10^{6}$ independent realizations of a Zeta distribution [^2] A discrete r.v. $W$ follows a Zeta distribution with parameter $\alpha$ if $\mathbb{P}\left(W=k\right)=\left(k^{\alpha}\zeta\left(\alpha\right)\right)^{-1}$ where $\zeta$ is the Riemann zeta function. The cdf of a Zeta distribution satisfies $\mathbb{P}(W\geq k)\sim k^{\alpha+1}/(\zeta(\alpha)(\alpha+1))$. This distribution is also known as Zipf’s distribution due to its relationship with Zipf’s law.   with parameter $\alpha=\beta-1$. (b) Behaviour of the Hill estimator for the same dataset.
@@
[…]
(a) $\widehat{\beta}_{n}(k)$ estimator
(b) Hill estimator
Figure 2: (a) Behaviour of $\widehat{\beta}_{n}(k)$ for different values of $k$, to estimate the parameter $\beta=0.85$ based on a dataset of $10^{6}$ independent realizations of the random variable $X=10^{-1}\lfloor 10U^{-1/\beta}\rfloor$. (b) Behaviour of the Hill estimator for the same dataset.
@@ 2.2 Refined ’Bias vs Variance’ Analysis - Examples
To illustrate this trade-off, we present the following Monte-Carlo experiment: We generate $10^{4}$ samples of a heavy-tailed distribution and calculate $\widehat{\beta}_{10^{4}}({k})$ for all admissible values of $k$, we repeat this experiment 100 times, and then we calculate the mean and the 95% confidence interval of $\widehat{\beta}_{10^{4}}({k})$ for each value of $k$. The results of these simulations, for the cases where $L(n)$ is asymptotically constant and $L(n)$ is logarithmic, are presented in Figures 3(a) and 3(b). As expected, the behaviour of the estimator is way better in the former case than in the latter.
@@ 3.1 Background and Preliminaries
The class of regenerative chains includes a wide variety of Markov processes, including all chains with countable state spaces (where any recurrent state serves as an accessible atom). This class also incorporates many Markov models frequently employed in operations research and queuing theory (see, for instance, [3]). Further examples of (regular) regenerative chains are presented in Examples 2 and 3 below.
@@
    There exists a slowly varying function $L_{A}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that: $\forall n\geq 1$,
    $\mathbb{P}_{A}\left(\tau_{A}\geq n\right)=L_{A}(n)\cdot n^{-\beta}.$ (16)
@@
The inference technique for the regularity index $\beta$ of the chain $X$ developed in subsection 3.3 is based on characterization $(ii)$: the parameter $\beta$ is the tail index of a discrete generalized Pareto r.v., the regeneration time namely, i.e. the conditional survivor function of $\tau_{A}$ given $X_{0}\in A$. Incidentally, notice that the parameter $\beta$ does not depend on the atom $A$ considered (in contrast to the estimator analysed in subsection 3.3). Based on a (random) number $N_{n}$ of (dependent) realizations of the regenerative time, namely
@@ 3.3 Regularity Index of a Regular Chain - Statistical Inference
provided that $\widehat{p}^{(N_{n})}_{k+1}>0$. We point out that, due to the randomness of $N_{n}$, (19) is a biased (strongly consistent) estimator of $p_{l}=\mathbb{P}_{A}(\tau_{A}>e^{l})$ for any $l\in\mathbb{R}$. The estimator (20) is of the same form as (2) except that the number $N_{n}$ of observations is random, and the observations are not independent anymore (in particular $S_{1}+\ldots+S_{N_{n}}\leq n$).
@@
Suppose that the atomic chain $X$ is $\beta$-regular with $\beta\in(0,1)$. Let $\nu$ be its initial probability distribution. If, as $n\rightarrow+\infty$, we $\mathbb{P}_{\nu}$ almost-surely have $k_{n}\to+\infty$ so that $(\ln n)\exp(k_{n}\beta)/n=o(L_{A}(\exp(k_{n}))$, the estimator (20) is strongly consistent:
@@
In particular, strong consistency holds for $\widehat{\beta}_{N_{n}}(\ln N_{n})$.
Notice that the choice of $k_{n}=\ln(N_{n})$ is universal and does not depend on the value $\beta\in(0,1)$.
In order to show the finite sample behaviour of our estimator in the Markovian scenario, we have simulated 100 independent trajectories of $10^{7}$ points for two Markov chains (a Simple Symmetric Random Walk where $\beta=0.5$ and a Bessel Random Walk with $\beta=0.6$), and we have computed $\widehat{\beta}_{N_{10^{7}}}(\ln N_{10^{7}})$ for each one of the trajectories. As a comparison, we also include the estimations using the statistic $\widetilde{\beta}_{10^{7}}$ described in (18). Figure 7 shows the KDE for both estimators.
{subcaptiongroup}
Figure 5: Simple symmetric random walk $\beta=0.5$
Figure 6: Bessel random walk with $\beta=0.6$
Figure 7: KDE for the estimators $\widehat{\beta}_{N_{10^{7}}}\left(\ln{N_{10^{7}}}\right)$ and $\widetilde{\beta}_{N_{10^{7}}}(A)$ in different $\beta$-regular chains.
@@
(On investigating convergence rates) Due to the impossibility of tightly controlling the sequence $N_{n}$ by a deterministic quantity in probability and the non-linearity of the estimator, we have not been able to extend to the Markovian case the asymptotic normality results of Theorem 2.2 and Corollary 2.2 via Anscombe’s theorem [19, Theorem 7.3.2]. Heuristically, if in Theorem 2.2 we take $k_{n}=\ln N_{n}$ and replace $N_{n}$ by its approximate expectation $n^{\beta}L_{\nu,A}(n)$ [24, Lemma 3.3], we would get a convergence rate of order $n^{-\beta\left(1-\beta\right)/2}L_{1}\left(n\right)$, where $L_{1}\left(n\right)$ is the slowly varying function given by $\sqrt{L_{\nu,A}(n^{\beta}L_{\nu,A}(n))/L_{\nu,A}(n)^{1-\beta}}$. This suggests a convergence rate of order $n^{-\beta(1-\beta)/2}$ when $L_{\nu,A}\sim C>0$. However, we have not been able to prove this claim.

### v3

@@ 1 Introduction
where $L:\mathbb{R}_{+}\to\mathbb{R}$ is a slowly varying function, i.e. such that $L(\lambda z)/L(z)\rightarrow+1$ as $z\rightarrow+\infty$ for any $\lambda>0$, see [7]. Such discrete power law probability distributions, also referred to as generalized Zipf’s laws sometimes, are often used to model the distribution of discrete data exhibiting a specific rank-frequency relationship, namely when the logarithm of the frequency and that of the rank order are nearly proportional. Such a phenomenon has been empirically observed in many ranking systems: in quantitative linguistics (i.e. when analysing word frequency law in natural language, see e.g. [30]) in the first place, as well as in a very wide variety of situations, too numerous to be exhaustively listed here. One may refer to [37], [26] or [40] among many others. In this paper, we first consider the issue of estimating the parameter $\beta$ involved in (1) (supposedly unknown, like the function $L$) in the classic (asymptotic) i.i.d. statistical setting, i.e. based on an increasing number $n\geq 1$ of independent copies $S_{1},\;\ldots,\;S_{n}$ of the generic r.v. $S$. Statistical inference for discrete heavy-tailed distributions has not received much attention in the literature and still poses methodological problems that this article seeks to resolve. Most of the very few dedicated methods documented either deal with very specific cases as in e.g. [15], [28] and [11] or else consist in empirically applying techniques originally designed for continuous heavy-tailed distributions to discrete data after first adding independent uniform noise [39, 21, 22].
@@
The vast majority of the regular variation index estimators proposed in the literature, Hill or Pickand estimators in particular [19, 35], are based on order statistics, which causes obvious difficulties in the discrete case because of the possible occurrence of many ties. Indeed, in that case, many spacings are equal to 00 which causes these estimators to behave erroneously. This is stressed in [28] (see their figures 1 and 2). More importantly, [28] constructs an example of a family of discrete value heavy-tailed distributions that do not satisfy any second-order condition so that the asymptotic normality of the Hill estimator cannot be proved directly. In contrast, the estimator under study here is based on the analysis of the probability of exponentially separated tail events. It simply rests on the fact that, as can be immediately deduced from (1), we have $\ln(p_{k})-\ln(p_{k+1})=\beta+\ln(L(e^{k})/L(e^{k+1}))$, where $\ln(x)$ denotes the natural logarithm of any real number $x>0$ and $p_{l}=\mathbb{P}(S>e^{l})$ for all $l\in\mathbb{N}$, and that $L(e^{k+1})/L(e^{k})$ is expected to be very close to $1$ for $k\in\mathbb{N}$ chosen sufficiently large. A natural (plug-in) inference technique can be then devised by replacing the tail probabilities $p_{l}$ with their empirical versions $\widehat{p}^{(n)}_{l}=(1/n)\sum_{i=1}^{n}\mathbb{I}\{S_{i}>e^{l}\}$ for $l\in\mathbb{N}$, where $\mathbb{I}\{\mathcal{A}\}$ means the indicator function of any event $\mathcal{A}$. This yields the estimator
@@
Although estimation of the parameter $\beta$ in (1) in the discrete i.i.d. setting is an important issue in itself, the present paper also finds its motivation in the problem of recovering statistically the regularity index of a regenerative regular $\beta$-null-recurrent Markov chain $X=(X_{n})_{n\in\mathbb{N}}$, based on the observation of a finite sample path $X_{1},\;\ldots,\;X_{n}$ with $n\geq 1$. As explained in [8, 9], for regular Markov chains, the regularity index $\beta$ controls the (sublinear) rate at which the number of visits to any given Harris set increases with observation time $n$, no matter the initial distribution. These Markov chains are strongly non-stationary and are not mixing in any sense. Typical examples of these chains are random walks and their variations, for instance Bessel random walks (see [2]). Such processes appear also naturally in many applications: for population dynamics or stochastic growth models in demography, see for instance [1] and the references therein; for branching processes, Galton-Watson processes with immigration (or population dependent) also exhibit a fat tail behaviour of the time returns of the chain, see [34, 41, 23].
@@
In the regenerative case (i.e. when the chain $X$ possesses an accessible atom, a Harris set on which the transition probability is constant), the distribution of the regenerative time, the return time to the atom, is a discrete generalized Pareto (1) and the parameter $\beta$ is its tail index. Due to the non-standard behaviour of traditional estimators in this context, statistical inference for null-recurrent Markov chains is very poorly documented in the literature [25, 24, 27, 17] and, to the best of our knowledge, estimation of the key quantity $\beta$ has not received much attention. It is also the goal of this article to extend the use of the estimator (2) to the case where the $S_{i}$’s are the successive durations between the consecutive regeneration times up to time $n$. The main difficulty naturally arises from the fact that the number $1+N_{n}\geq 0$ of regeneration times (and thus the number of durations) is now random, and it is not independent of the variables $S_{1},\ldots,S_{N_{n}}$ (in particular, $S_{1}+\cdots+S_{N_{n}}\leq n$ by construction). We show that the estimator remains strongly consistent in this scenario and we provide non-asymptotic upper confidence bounds for the absolute deviations between $\widehat{{\beta}}_{N_{n}}\left(k\right)$ and $\beta$. For illustration purposes, numerical experiments have been carried out, providing empirical evidence of the relevance of the estimation method promoted. Extension to the general case of (pseudo-regenerative) null-recurrent chains is also discussed, the difficulties inherent in applying the methodology originally proposed in [6] in the positive recurrent case to mimic regenerative Nummelin extensions [33] being explained at length.
@@
The paper is organized as follows. A thorough analysis of the behaviour of the estimator (2) in the i.i.d. case, illustrated by numerical experiments, is first carried out in section 2. Some of the results thus established (consistency and non-asymptotic bounds) are next extended in section 3 to the regenerative regular Markovian setup, when the estimator is computed based on a single finite-length trajectory of the atomic chain. Experimental results are also displayed and the main barrier to the extension of the methodology promoted to general (i.e. pseudo-regenerative) regular null-recurrent chains is also discussed therein. Technical proofs are deferred to the Appendix section.
@@ 2.1 Main Results - Confidence Bounds and Limit Theorems
Suppose that, as $n\rightarrow+\infty$, we have $k_{n}\to+\infty$ and $(\ln n)\exp(k_{n}\beta)/n=o(L(\exp(k_{n}))$. Then, we have:
@@
In particular, as stated below, strong consistency is guaranteed when $k_{n}$ is of logarithmic order.
@@
Of course, the condition (5) on $k_{n}$ can be difficult to check in practice. This is a common issue in tail estimation and in the statistical analysis of extreme values more generally. The choice of the hyperparameter $k$ somehow governs the (asymptotic) bias-variance trade-off: the estimator (2) is expected to have a large variance when $k$ is large and to have a large bias if $k$ is too small. As depicted in Fig. 1(a), to choose $k$, one may use the same approach as that originally proposed for the Hill estimator (see e.g. [36]), which consists of plotting the values of (2) for a range of values of $k$ and choosing $k$ in a region where the estimator shows some stability. Section 3.2 of [10] suggests an adaptive algorithm for selecting $k$ for $\widehat{\beta}_{n}$, assuming the stricter condition that $F$ belongs to the class of second-order Pareto distributions.
We point out that, contrary to (variants of) the Hill inference method, the integer hyperparameter $k$ does not refer here to the number of order statistics involved in the estimator. Instead, it determines a threshold that defines the tail probability we want to estimate. This implies that the range of $k$ is the set of integers between $1$ and $\max_{1\leq i\leq n}\ln S_{i}-1$, while, for classical tail index estimators, the hyperparameter varies between $1$ and $n$.
@@
Figure 1: (a) Behaviour of $\widehat{\beta}_{n}(k)$ for different values of $k$, to estimate the parameter $\beta=0.8$ based on a dataset of $10^{6}$ independent realizations of a Zeta distribution [^2] A discrete r.v. $W$ follows a Zeta distribution with parameter $\beta$ if $\mathbb{P}\left(W=k\right)=1/(k^{\beta-1}\zeta(\beta-1))$ where $\zeta$ is the Riemann zeta function. Its cdf satisfies $\mathbb{P}(W\geq k)\sim k^{\beta}/(\beta\zeta(\beta-1))$. This distribution is also known as Zipf’s distribution due to its relationship with Zipf’s law. . (b) Behaviour of the Hill estimator for the same dataset. Notice that the range of the x axis in both graphs is different, in the case of $\widehat{\beta}_{n}(k$ the range of $k$ is the set of integers between $1$ and $\max_{1\leq i\leq n}\ln S_{i}-1$ while in the case of the Hill method, the x-axis is the number of order statistics involved in the estimator, hence, it goes from 1 to $n$.
Remark 1.
(Quantiles vs tail probabilities) In Extreme Value Theory, it is customary to estimate the tail index by looking at the largest values, i.e. a fraction of the order statistics. The general idea consists of fixing the tail probabilities and then using the order statistics (i.e. empirical quantiles) to estimate the quantiles [12, Chapter 3]. The estimator proposed here follows the converse approach: we fix the thresholds/quantiles, next use the data to estimate the tail probabilities and finally use these estimations in order to mimic the limit behavior of (3). In general, estimators based on order statistics are difficult to analyse in a non-asymptotic way, whereas our estimate is based on probabilities of well-chosen tail events.
@@
This shows that if $k_{n}$ satisfies the conditions of Theorem 2.1 and $ne^{-k_{n}(\beta+2)}$ goes to 00, then the conditions of Corollary 2.2 are satisfied and $\sqrt{np_{k_{n}}}({\widehat{\beta}_{n}({{k_{n}}})-\beta})$ converges in distribution to a centered normal random variable. Moreover, if we take $k_{n}=A\ln n$ (as in Corollary 2.1) then, the asymptotic normality is guaranteed as long as $1/(\beta+2)<A<1/\beta$.
@@ 2.2 Refined ’Bias vs Variance’ Analysis - Examples
To illustrate this trade-off, we present the following Monte-Carlo experiment: We generate $10^{4}$ samples of a heavy-tailed distribution and calculate $\widehat{\beta}_{10^{4}}({k})$ for all admissible values of $k$, we repeat this experiment 100 times, and then we calculate the mean and the 95% confidence interval of $\widehat{\beta}_{10^{4}}({k})$ for each value of $k$. The results of these simulations, for the cases where $L(n)$ is asymptotically constant and $L(n)$ is logarithmic, are presented in Figures 2(a) and 2(b). As expected, the behaviour of the estimator is way better in the former case than in the latter.
@@ 3.1 Background and Preliminaries
The class of regenerative chains includes a wide variety of Markov processes, including all chains with countable state spaces (where any recurrent state serves as an accessible atom). This class also incorporates many Markov models frequently employed in operations research and queueing theory (see, for example, [3]). Further examples of (regular) regenerative chains are presented in Examples 2 and 3 below.
@@
    There exists a slowly varying function $L_{A}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that: $\forall m\geq 1$,
    $\mathbb{P}_{A}\left(\tau_{A}\geq m\right)=L_{A}(m)\cdot m^{-\beta}.$ (16)
@@
The inference technique for the regularity index $\beta$ of the chain $X$ developed in subsection 3.3 is based on characterization $(ii)$: the parameter $\beta$ is the tail index of a discrete generalized Pareto r.v., the regeneration time namely, i.e. the conditional survivor function of $\tau_{A}$ given $X_{0}\in A$. Incidentally, notice that the parameter $\beta$ does not depend on the atom $A$ considered (in contrast to the estimator analysed in subsection 3.3).
Based on a (random) number $N_{n}$ of (dependent) realizations of the regenerative time, namely
@@ 3.3 Regularity Index of a Regular Chain - Statistical Inference
provided that $\widehat{p}^{(N_{n})}_{k+1}>0$. We point out that, due to the randomness of $N_{n}$, (19) is a biased (strongly consistent) estimator of $p_{l}=\mathbb{P}_{A}(\tau_{A}>e^{l})$ for any $l\in\mathbb{R}$. The estimator (20) is of the same form as (2) except that the number $N_{n}$ of observations is random, and it is not independent of the sequence $\{S_{i}\}$ of observations (in particular $S_{1}+\ldots+S_{N_{n}}\leq n$).
Obtaining properties of this new estimator requires understanding the behavior of $N_{n}$, to this, the key is the following result, proved in [25], which gives the asymptotic distribution of $N_{n}$.
Proposition 3.3.
([25], Theorem 3.2) If $X$ is an atomic $\beta$-regular Markov chain with initial probability distribution $\nu$, then
[…]
where $M_{\beta}(1)$ is a Mittag-Leffler distribution with parameter $\beta$.
From Proposition 3.3 we can deduce the following corollary which provides a deterministic control of $N_{n}$ with high probability.
Corollary 3.1.
For any $\delta\in(0,1)$, let $H_{\delta}>R_{\delta}>0$ be such that
$\mathbb{P}\left(\frac{M_{\beta}(1)}{\Gamma(1-\beta)}\in\left[R_{\delta},H_{\delta}\right]\right)\geq 1-\delta/2.$
Then, there exists $n_{\delta}$ such that
$\forall n\geq n_{\delta}\quad\mathbb{P}_{\nu}\left(N_{n}\in\left[\frac{R_{\delta}}{\overline{F}(n)},\frac{H_{\delta}}{\overline{F}(n)}\right]\right)>1-\delta.$
Equipped with Corollary 3.1, we can provide the following result, which is the analogue of Proposition 2.1 in the Markovian case.
Proposition 3.4.
[…]
$\left|\widehat{\beta}_{N_{n}}(k)-\beta\right|\leq\frac{60H_{\delta}}{R_{\delta}}\sqrt{\frac{v_{n}(\delta)}{p_{k+1}}}+\left|\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right)\right|.$
@@
Suppose that the atomic chain $X$ is $\beta$-regular with $\beta\in(0,1)$. Let $\nu$ be its initial probability distribution. If $k_{n}\to+\infty$ s.t. $(\ln n)\exp(k_{n}\beta)/n=o(L_{A}(\exp k_{n}))$ as $n\rightarrow+\infty$, then the estimator (20) is strongly consistent:
@@
In particular, strong consistency holds for $\widehat{\beta}_{N_{n}}(A\ln N_{n})$ with $A<1$.
@@
(On investigating convergence rates) Due to the impossibility of tightly controlling the sequence $N_{n}$ by a deterministic quantity in probability and the non-linearity of the estimator, we have not been able to extend to the Markovian case the asymptotic normality results of Theorem 2.2 and Corollary 2.2 via Anscombe’s theorem [18, Theorem 7.3.2]. Heuristically, if in Theorem 2.2 we take $k_{n}=\ln N_{n}$ and replace $N_{n}$ by its approximate expectation $n^{\beta}L_{\nu,A}(n)$ [25, Lemma 3.3], we would get a convergence rate of order $n^{-\beta\left(1-\beta\right)/2}L_{1}\left(n\right)$, where $L_{1}\left(n\right)$ is the slowly varying function given by $\sqrt{L_{\nu,A}(n^{\beta}L_{\nu,A}(n))/L_{\nu,A}(n)^{1-\beta}}$. This suggests a convergence rate of order $n^{-\beta(1-\beta)/2}$ when $L_{\nu,A}\sim C>0$. However, we have not been able to prove this claim.
@@
3.4 Simulation Experiments
In order to analyze empirically the finite sample behavior of our estimator in the Markovian case, we dedicate this subsection to a simulation example. As a generative model for our experiments we use the Bessel random walks, defined in Example 2, as it will allow us to study the accuracy of our estimator, not just for different sample sizes, but also for different values of $\beta$.
For different values of $\beta$ and $n$ we have performed the following two-fold experiment:
  1. 1.
    First, we have generated one path of a $\beta$-regular Bessel random walk of length $n$. We have then computed the estimator $\widehat{\beta}_{N_{n}}(k)$ based on this trajectory and plotted the results in order to determine the value $k_{\beta,n}$ where it first stabilizes.
  2. 2.
    We have simulated $2000$ independent paths of a $\beta$-regular Bessel random walk of length $n$. Then, we have computed the estimator $\widehat{\beta}_{N_{n}}(k_{\beta,n})$ based on each of these trajectories, as well as the estimator $\widetilde{\beta}_{n}(\{0\})$.
The results of this experiment are presented in Table 1. They show that the accuracy of our estimator greatly increases as the length $n$ gets larger (notice that the bias and variance are both divided by two in order of magnitude when $n$ increases from $10^{3}$ to $10^{6}$). The histograms shown in Figure 6 hint a possible asymptotic normality of the estimator $\widehat{\beta}_{N_{n}}$. The simulation study also suggests that for moderately large sample sizes ($n\geq 10^{5}$) our estimator $\widehat{\beta}_{N_{n}}$ outperforms $\widetilde{\beta}_{n}$ as it has less bias and comparable variance, while for smaller sample sizes $\widetilde{\beta}_{n}$ works better. This behavior can be attributed to the nature of our estimator, which focuses on estimating the probabilities of tail events $\{S>e^{k}\}$ to approximate the quantity $\beta(k)$ and use the fact that, as $k$ tends to infinity, $\beta(k)$ converges to $\beta$ (see (3)). When $n$ is small, there are fewer sample points available in the tail, which forces us to select smaller values of $k$ to obtain accurate estimations of the tail probability, which worsens the estimation of $\beta$.
[…]
$\widehat{\beta}_{N_{n}}$ $\widetilde{\beta}_{n}(\{0\}$ $\widehat{\beta}_{N_{n}}$ $\widetilde{\beta}_{n}(\{0\})$
0.5 $10^{3}$ 2 0.22255 0.07987 0.11062 0.01932
$10^{4}$ 2 0.12246 0.06697 0.06032 0.01317
$10^{5}$ 3 0.01024 0.05424 0.02921 0.00825
$10^{6}$ 3 0.00343 0.04622 0.01309 0.00591
0.7 $10^{3}$ 2 0.20290 0.13154 0.12575 0.01580
$10^{4}$ 3 0.02885 0.10598 0.05673 0.00969
$10^{5}$ 3 0.00500 0.08249 0.01680 0.00604
$10^{6}$ 3 0.00476 0.06961 0.00389 0.00411
0.9 $10^{3}$ 3 0.27182 0.20158 0.15869 0.00966
$10^{4}$ 3 0.03128 0.16307 0.04685 0.00539
$10^{5}$ 4 0.02338 0.13529 0.01807 0.00286
$10^{6}$ 5 0.00704 0.11502 0.00606 0.00177
Table 1: Bias-Variance results of the simulation example.
{subcaptiongroup}
Figure 4: $\beta=0.9$, $n=10^{4}$
Figure 5: $\beta=0.5$, $n=10^{6}$
Figure 6: KDE for the estimators $\widehat{\beta}_{N_{n}}$ and $\widetilde{\beta}_{N_{n}}(\{0\})$.